Fly-by-Night Airlines has a peculiar rule about luggage: The length l and width w of a bag must add up to at most 57 inches, and the width w and height h must also add up to at most 57 inches. What are the dimensions of the bag with the largest volume that Fly-by-Night will accept?
Answer #1

Maximize LxHxW

Constraint L+W=57 and W+H=57

in this case L must equal H since the sums are both 57

So the maximization becomes L2xW

the constraint is L+W=57

Solving the constraint for W yields

W=57-L

place this into the maximization eqn

f(L)=L^{2}(57-L)= -L^{3}+57L^{2}

now we take the derivative of this equation and determine the critical points (where the derivative=0)

derivative

f'(L)=-3L^{2}+114L=0

use the quadratic equation to solve

this yields L=0 and L=38

obviously L cannot = 0

solve for the other variables

if L=38 then H=38

L+W=57

W=57-L=57-38=19

L=38

H=38

W=19

Vmax=38*38*19=27436 cubic inches
Answer #2

The lenth l and width w of a bag must add up to at most 57 inches.

\Rightarrow l+w\le 57  \Rightarrow l\le 57-w...(1)

The height h and width w of a bag must add up to at most 57 inches.

\Rightarrow h+w\le 57  \Rightarrow h\le 57-w...(2)

Now, we have to maximize the volume, lwh, of such a bag.

lwh\le (57-w)w(57-w) [ Using (1), (2)]

\Rightarrow lwh\le w(57-w)^2=w(w^2-114w+3249)=w^3-114w^2+3249w

Now, we have to maximize g(w)=w^3-114w^2+3249w to get the answer.

g'(w)=3w^2-2\times114w+3249=3(w^2-76w+1083)=3(w-57)(w-19)

Now, when, g(w) attains extremum, then, g'(w)=0\Rightarrow 3(x-19)(x-57)=0\Rightarrow w=19,57

g''(w)=3(2w-76)

g attains maximum when g''(w)<0.

We check that, g''(19)=3(2\times 19-76)=-3\times 38=-114<0

and, g''(57)=3(2\times 57-76)=3\times 38=114>0

\Rightarrow g attains maximum when w=19

\Rightarrow lwh\le 19^3-114\times 19^2+3249\times 19=27436

The dimensions of the bag with largest volume that fly-by-night will accept are

l=57-19=38 inches, w=19 inches, h=57-19=38 inches.